nhihoaa:

a bag contains quarters and dimes there are 15 coins in the bag. the value of the coins is $2.7. determine the number of quarters and dimes are in the bag

Vocaloid:

gotta write a system of equations for this let's let q = # of quarters and d = # of dimes we have 15 coins, so q + d = 15 0.10 * d gives the monetary value of the dimes since each dime is worth 10 cents 0.25 * q gives the monetary value of the quarters since each quarter is worth 25 cents so adding up the values, 0.10d + 0.25q = 2.7 solve the system q + d = 15 0.10 + 0.25q = 2.7